# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v24

@@ 3.8 Named Entity Recognition Tools
The table VI provides a comparative summary of the discussion above. It should be noted that all three of these models only tag NEs of types: person names, location names and organization names. The Ananya system by Manamini et al. [118] is available to download at GitHub 106106106http://bit.ly/2XrwCoK. The data and code for the approaches by Dahanayaka and Weerasinghe [222] and by Senevirathne et al. [223] are not accessible to the public. Azeez and Ranathunga [224] proposed a fine-grained NER model for Sinhala building on their earlier work on NER [118] and PoS tagging [117]Anuruddha [225] proposed a method based on reinforcement learning for Sinhala NER. A Sinhala NER system restricted to the sports domain was introduced by Wijesinghe and Tissera [226], where they attempted to utilise classical machine learning models. The work by Mallikarachchi et al. [227] used support vector machines to detect Sinhala named entities of types person, location and, organization. Even though the main task reported in the study by [228] is clustering Sinhala news articles, they used NER tools to do so. The work by Gurgurov et al. [229] compared the performance of LLMs for the Sinhala NER task. The highest results are obtained for this task in the Masked Language Modeling objective for Sinhala. They have used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA). A data set for Sinhala NER named multiNER107107107https://github.com/suralk/multiNER was introduced by Ranathunga et al. [168]. It includes parallel NER data for Sinhala, Tamil, and English.
@@ 3.10 Text Classification
Text classification is a popular application on the semantic layer of the NLP stack. A very basic Sinhala text classification using Naïve Bayes Classifier, Zipf’s Law Behavior, and SVMs was attempted by Gallege [235]. A smaller implementation of Sinhala news classification has been attempted by de Silva [28]. As mentioned in Section 3.2, their news corpus is publicly available108108108https://osf.io/tdb84/. Another attempt at Sinhala text classification using six popular rule-based algorithms was done by Lakmali and Haddela [236]. Even though they talk about building a corpus named SinNG5, they do not indicate of means for others to obtain the said corpus. Another study by Kumari and Haddela [237] utilizes the SinNG5 corpus as the data set for their attempt to use LIME [238] for human interpretability of Sinhala document classification. However, they too do not provide access to the corpus. Nanayakkara and Ranathunga [239] have implemented a system which uses corpus-based similarity measures for Sinhala text classification.  Gunasekara and Haddela [240, 241] claim to have created a context-aware stop word extraction method for Sinhala text classification based on simple TF-IDF. An LSTM based textual entailment system for Sinhala was proposed by Jayasinghe and Sirts [242]Demotte and Ranathunga [243] proposed a dual-state capsule network architecture for text classification where they demonstrated their methodology on the Sinhala data set established by Senevirathne et al. [244]. The attempt by Sameemdeen and Selvanthan [245] considers three classical machine learning algorithms (Naïve Bayes,SVM,and KNN) and then goes on to briefly discusses the pros and cons of previous attempts by: Nanayakkara and Ranathunga [239]Gunasekara and Haddela [240]Lakmali and Haddela [236], and Buddhika et al. [246]. Then they propose active learning [247, 248] as an alternative. However, no experimental results of how active learning would improve Sinhala text classification are given. The study by Bandara et al. [249] proposed and ontology-based approach for Sinhala fake news detection. However, their literature survey did not cover seminal papers in OBIE such as the work by Wimalasuriya and Dou [37]. This has impacted the overall methodology that was presented. A novel, domain-idependant, and domain-adaptive text classification framework named AdaptText for Sinhala text classification was proposed by Kodithuwakku and Hettiarachchi [250]. A simple TF-IDF based Sinhala text classification system was proposed by Koralage [251]. The study by [252] used Genetic Algorithm on Lucene109109109https://lucene.apache.org/ search queries to obtain interpretable classification models for Sinhala text documents. The approach proposed by Rathnayake et al. [253] uses adapter-based [254, 255, 256, 257, 258, 259] fine-tuning on XLM-R [260], for code-mixed and code-switched text. Kirindage and Godewithana [261] used LDA [262] to develop Sinhala news topic hierarchies and categorize Sinhala news documents using the said topic hierarchies. Hettigoda [263] collected English-Sinhala code-mixed comments from Facebook pages belonging to Clothing industrial online businesses in Sri Lanka and classified them into the classes: Inquiries, Maybe Inquiries, and Not Inquiries. A comparative analysis on BERT based models for Sinhala text classification was conducted by Dhananjaya et al. [107]. Both their code110110110https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus1111111113https://tinyurl.com/42un7a9y, Sinhala News source classification data set112112112https://bit.ly/3PBTyfo, Sinhala News category classification data set113113113https://bit.ly/3AsfK7u, and Sinhala writing style classification data set114114114https://bit.ly/3PARuUS) are publicly available. Faumi et al. [136] also worked on Sinhala writing style identification (author identification) using their own data set115115115https://github.com/cipherdragon/SimpleAA which includes Romanized (transliterated) Sinhala text.
The study by Wijayarathna and Jayalal [264] used classical machine learning techniques including Random Forest to classify fake news in Sinhala on Twitter. Chathuranga and Ranathunga [265] used capsule-based methods recommended by Senevirathne et al. [244] to classify Sinhala-English code-mixed data. The work by Weerasiri et al. [266] compared Word2Vec, FastText and Doc2vec [267] against TF-IDF for Sinhala news document classification. The study by Caldera et al. [268] used stacked LSTM to classify Sinhala and Singlish text discussing COVID-19 on Youtube and Twitter.
@@ 3.11 Sentiment Analysis
A simple MLP-based method to classify sentiments in Sinhala text was initially proposed by Medagoda [269] based on their prior work [270]. A word2vec based tool116116116http://bit.ly/2QKI9Np for sentiment analysis of Sinhala news comments is available. A methodology for constructing a sentiment lexicon for Sinhala Language in a semi-automated manner based on a given corpus was proposed by Chathuranga et al. [271]Demotte et al. [272] proposed a sentiment analysis system117117117https://github.com/theisuru/sentiment-tagger based on sentence-state LSTM Networks for Sinhala news comments. In the subsequent work [273], they used word similarity to generate a Sinhala semantic lexicon. They followed this up with a further study [244] which discussed a number of other deep learning models such as RNN and Bi-LSTM in the domain of Sinhala sentiment analysis. Jayasuriya et al. [274] proposed a method to classify Sinhala posts in the domain of sports into positive and negative class sentiments. Ranathunga and Liyanage [275] claimed that using word embedding models as semantic features can compensate for the lack of well developed language-specific linguistic or language resources in the case of analysing sentiment of Sinhala news comments. Jayasuriya et al. [276] conducted a comparative study between word N-grams and character N-grams in the task of semantic classification of Sinhala content in social media. Which they soon followed up with an ensemble approach [277]. The work by Karunarathne [278] used word embedding to analyse the sentiment of manually annotated Sinhala Tweets. Abeyratne and Jayaratne [279] conducted a multi model analysis on classifying Sinhala songs by emotion.
@@
The work by Jayawickrama et al. [280, 281] used the data set released by Wijeratne and de Silva [99] to predict the reactions induced by Sinhala Facebook posts. They then extended the work [282] and compared the results obtained with their data set against that of Senevirathne et al. [244]Aththanayaka and Naleer [283] used Random forest, Support vector machines, and Multinomial Naïve Bayes models to analyse sentiment in Sinhala-English code-mixed text from social media. The ACTSEA dataset118118118https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al. [135]. The dataset contains 318,308 Sinhala tweets annotated with emotions. The study by Dhananjaya et al. [284] uses a sentiment lexicon of a high‐resource language to fine-tune Pre‐trained multilingual language models (PMLMs) such as mBERT and XLM‐R on an intermediate task which in turn is then used on the sentiment classification task of Sinhala. A sentiment-annotated dataset of Sinhala-English code mixed YouTube comments has been created by Uthpala and Thirukumaran [285]. However, given that even their paper is behind a paywall, there is no free and public way to access this dataset. The work by Gurgurov et al. [229] compared the performance of LLMs for the Sinhala sentiment analysis task. They report the highest results for the Masked Language Modeling objective for Sinhala in both of their configurations where they used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA). Liyanaarachchi et al. [286] used character-level embedding to classify sentiments of Sinhala news comments as positive or negative. They claim that the 1-of-m embedding outperforms Log-m embedding for this task. Bandaranayake and Usoof [287] used BERT, DistilBERT, RoBERTa, and XLM-R to analyse the sentiment of Sinhala news comments [244]. They report that XLMR-large model achieves the best result. Fernando and Ranathunga [288] in their study on improving cross-lingual representation of multilingual language models for low-resource languages using linguistic entity masking used Sinhala code mixed sentiment analysis as one of the evaluations.
@@ 3.12 Hate Speech Detection
A machine learning approach to detect hate speech in Sinhala was proposed by De Silva [290]. A feature model and a data set119119119https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [128]Sandaruwan et al. [291] have attempted to identify abusive Sinhala comments in social media using text mining and machine learning techniques. A cyberbullying comment classification study for Sinhala was conducted by [292] where they used classical machine learning algorithms. The study by Hettiarachchi et al. [293] used classical machine learning methods to detect hate speech in Romanized Sinhala social media posts. While the basic idea is the same, they have avoided mentioning transliteration in their paper. The study by Samarasinghe et al. [294] proposed using CNNs for detecting hate speech in Sinhala text. Kariyawasam [295] proposed a machine learning approach for identifying toxic Sinhala language on social media. Guruge et al. [296] used an ensemble of Naïve Bayes, Support Vector Machine, XGBoost, MLP, and AdaBoost to detect hate speech in 49019 Tweets they collected from February of 2021 to April of 2021. Sandaruwan et al. [297] collected 3000 Sinhala comments corpus and conducted Multinomial Naïve Bayes to detect hate speech. Sheran [298] claims to have used machine learning to detect hate speech written in Sinhala or Singlish on social media. A deep learning approach for detecting hate speech in Sinhala tweets was explored by Munasinghe and Thayasivam [299]. The study by Shalinda and Munasinghe [300] utilized classical machine learning techniques such as Linear Support vector classifier, Random Forest Classification, SGD classifier, Logistic Regression, XGBoost classifier and multinomial Naive Bayes classifier on both Sinhala and Singlish (Romanized Sinhala) to identify hate words. They report that the SGD classifier over TF-IDF with uni-grams and bi-grams gives them the highest accuracy. Gamage et al. [301] conducted a comparative analysis on a number of embedding systems as well as classical frequency-based methods for Sinhala hate speech detection. The study by Fernando et al. [302] also claims to use machine learning and deep learning to detect hate speech in Sinhala. The study by Perera et al. [303] predicted Sinhala hate speech using user behaviour on twitter by applying ensemble Machine Learning to classify them by the five big personality traits. The follow-up study by Perera et al. [129] analyses Sinhala hate speech propagation on Twitter. They have also published a dataset120120120https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets. Rajapaksha et al. [304] used deep learning to identify trending periods of hate topics on Twitter in Sinhala. Arachchi et al. [305] used a web-based tool to translate between Sinhala and English to detect inappropriate word usage in Sinhala. The study by Ehelepola [306] analyzed Sinhala text from social media and e-commerce sites for hate speech. The work by Dikwatta et al. [307] used supervised algorithms to detect Sinhala hate speech in text from image posts on the Internet (memes). Wickramaarachchi et al. [308] used LSTM on BART to compare the title and description of Sinhala youtube videos to their audio in order to determine if they contain hate speech. The study by Muthuthanthri and Smith [309] used an annotated dataset from Facebook with TF-IDF and fastText embedding features with SVM, CNN, RNN, LSTM as well as BERT and GPT2. They concluded that BERT obtains the best result for detecting hate speech in English-Sinhala code-mixed data. Chavinda and Thayasivam [172] have used Multilingual Large Language Models (MLLMs) with Dual Contrastive Learning (DCL) to detect Sinhala hate speech on Facebook121121121https://huggingface.co/datasets/krishan-CSE/Facebook_Sinhala_Hate_Speech and Twitter122122122https://huggingface.co/datasets/krishan-CSE/Twitter_Sinhala_Hate_Speech data sets. They report TwHIN-BERT-base [310] model to have obtained the best results.
@@
After the release of the SOLD dataset by Ranasinghe et al. [140] for the hate speech detection task in Sinhala, a number of subsequent works have been conducted using it. Fernando and Deng [311] used traditional machine learning algorithms to predict Sinhala hate speech. Ranasinghe and Zampieri [312] compared the results of pre-trained mT5 on Sinhala tweets from the SOLD dataset against results for German, Spanish, Hindi, Korean, and Marathi. The work by Bestgen [313] compared the result of classical machine learning algorithms on simple n-gram features for the hate speech detection task on the SOLD dataset for Sinhala against results for Assamese and Bengali. Ranasinghe et al. [314] used the SOLD dataset for the Hate Speech And Offensive Content (HASOC) identification task123123123https://hasocfire.github.io/hasoc/2023/task1.html they introduced. In completing the task, the study by Narayan et al. [315] compared the effectiveness of a number of pre-trained deep learning models for Sinhala hate speech detection. They showed that XLM Roberta Base achieves a macro-f1 of 83.49% over the LSTM baseline of 75.30%. For the same task, Ojo et al. [316] used mBERT to identify hate and offensive content in Sinhala. The work by Rostamkhani and Eetemadi [317] compared the zero-shot performance of ChatGPT on the SOLD dataset against a number of other Sinhala models. They conclude that Zero-shot ChatGPT perform better in Sinhala hate speech detection than models that were not fine-tuned on SOLD. Dmonte et al. [141] used the SOLD data set to verify the new MT-Offense124124124https://github.com/LanguageTechnologyLab/MT-Offense data set they created by using NMT to translate the English offensive language benchmark dataset, OLID [142]. They report the best F1 of 0.550 for the translation done using nllb-200-3.3B [134].

### v25

@@ 3.8 Named Entity Recognition Tools
The table VI provides a comparative summary of the discussion above. It should be noted that all three of these models only tag NEs of types: person names, location names and organization names. The Ananya system by Manamini et al. [118] is available to download at GitHub 108108108http://bit.ly/2XrwCoK. The data and code for the approaches by Dahanayaka and Weerasinghe [223] and by Senevirathne et al. [224] are not accessible to the public. Azeez and Ranathunga [225] proposed a fine-grained NER model for Sinhala building on their earlier work on NER [118] and PoS tagging [117]Anuruddha [226] proposed a method based on reinforcement learning for Sinhala NER. A Sinhala NER system restricted to the sports domain was introduced by Wijesinghe and Tissera [227], where they attempted to utilise classical machine learning models. The work by Mallikarachchi et al. [228] used support vector machines to detect Sinhala named entities of types person, location and, organization. Even though the main task reported in the study by [229] is clustering Sinhala news articles, they used NER tools to do so. The work by Gurgurov et al. [230] compared the performance of LLMs for the Sinhala NER task. The highest results are obtained for this task in the Masked Language Modeling objective for Sinhala. They have used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA). A data set for Sinhala NER named multiNER109109109https://github.com/suralk/multiNER was introduced by Ranathunga et al. [168]. It includes parallel NER data for Sinhala, Tamil, and English. Rathnayake and Asanka [231] simply uses LLaMA for Sinhala NER.
@@ 3.10 Text Classification
Text classification is a popular application on the semantic layer of the NLP stack. A very basic Sinhala text classification using Naïve Bayes Classifier, Zipf’s Law Behavior, and SVMs was attempted by Gallege [237]. A smaller implementation of Sinhala news classification has been attempted by de Silva [28]. As mentioned in Section 3.2, their news corpus is publicly available110110110https://osf.io/tdb84/. Another attempt at Sinhala text classification using six popular rule-based algorithms was done by Lakmali and Haddela [238]. Even though they talk about building a corpus named SinNG5, they do not indicate of means for others to obtain the said corpus. Another study by Kumari and Haddela [239] utilizes the SinNG5 corpus as the data set for their attempt to use LIME [240] for human interpretability of Sinhala document classification. However, they too do not provide access to the corpus. Nanayakkara and Ranathunga [241] have implemented a system which uses corpus-based similarity measures for Sinhala text classification.  Gunasekara and Haddela [242, 243] claim to have created a context-aware stop word extraction method for Sinhala text classification based on simple TF-IDF. An LSTM based textual entailment system for Sinhala was proposed by Jayasinghe and Sirts [244]Demotte and Ranathunga [245] proposed a dual-state capsule network architecture for text classification where they demonstrated their methodology on the Sinhala data set established by Senevirathne et al. [246]. The attempt by Sameemdeen and Selvanthan [247] considers three classical machine learning algorithms (Naïve Bayes,SVM,and KNN) and then goes on to briefly discusses the pros and cons of previous attempts by: Nanayakkara and Ranathunga [241]Gunasekara and Haddela [242]Lakmali and Haddela [238], and Buddhika et al. [248]. Then they propose active learning [249, 250] as an alternative. However, no experimental results of how active learning would improve Sinhala text classification are given. The study by Bandara et al. [251] proposed and ontology-based approach for Sinhala fake news detection. However, their literature survey did not cover seminal papers in OBIE such as the work by Wimalasuriya and Dou [37]. This has impacted the overall methodology that was presented. A novel, domain-idependant, and domain-adaptive text classification framework named AdaptText for Sinhala text classification was proposed by Kodithuwakku and Hettiarachchi [252]. A simple TF-IDF based Sinhala text classification system was proposed by Koralage [253]. The study by [254] used Genetic Algorithm on Lucene111111111https://lucene.apache.org/ search queries to obtain interpretable classification models for Sinhala text documents. The approach proposed by Rathnayake et al. [255] uses adapter-based [256, 257, 258, 259, 260, 261] fine-tuning on XLM-R [262], for code-mixed and code-switched text. Kirindage and Godewithana [263] used LDA [264] to develop Sinhala news topic hierarchies and categorize Sinhala news documents using the said topic hierarchies. Hettigoda [265] collected English-Sinhala code-mixed comments from Facebook pages belonging to Clothing industrial online businesses in Sri Lanka and classified them into the classes: Inquiries, Maybe Inquiries, and Not Inquiries. A comparative analysis on BERT based models for Sinhala text classification was conducted by Dhananjaya et al. [107]. Both their code112112112https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus1131131133https://tinyurl.com/42un7a9y, Sinhala News source classification data set114114114https://bit.ly/3PBTyfo, Sinhala News category classification data set115115115https://bit.ly/3AsfK7u, and Sinhala writing style classification data set116116116https://bit.ly/3PARuUS) are publicly available. Faumi et al. [136] also worked on Sinhala writing style identification (author identification) using their own data set117117117https://github.com/cipherdragon/SimpleAA which includes Romanized (transliterated) Sinhala text. The study by Wijayarathna and Jayalal [266] used classical machine learning techniques including Random Forest to classify fake news in Sinhala on Twitter. Chathuranga and Ranathunga [267] used capsule-based methods recommended by Senevirathne et al. [246] to classify Sinhala-English code-mixed data. The work by Weerasiri et al. [268] compared Word2Vec, FastText and Doc2vec [269] against TF-IDF for Sinhala news document classification. The study by Caldera et al. [270] used stacked LSTM to classify Sinhala and Singlish text discussing COVID-19 on Youtube and Twitter. The work by Gunawardana and Thamarasee [271] attempted to predict the subsequent popularity of Sinhala language YouTube videos by analysing a number of associated factors. Ilangeshwaran et al. [272] introduced Saliency-based token Swap (SSwap), a data augmentation technique to improve Sinhala text classification. The work by Rathnayake and Asanka [231] uses LLaMA for Named entity classification in Sinhala documents and then runs Fuzzy C-Means clustering on it. Next, they utilise LLaMA again to propose names to the created clusters.
@@ 3.11 Sentiment Analysis
A simple MLP-based method to classify sentiments in Sinhala text was initially proposed by Medagoda [273] based on their prior work [274]. A word2vec based tool118118118http://bit.ly/2QKI9Np for sentiment analysis of Sinhala news comments is available. A methodology for constructing a sentiment lexicon for Sinhala Language in a semi-automated manner based on a given corpus was proposed by Chathuranga et al. [275]Demotte et al. [276] proposed a sentiment analysis system119119119https://github.com/theisuru/sentiment-tagger based on sentence-state LSTM Networks for Sinhala news comments. In the subsequent work [277], they used word similarity to generate a Sinhala semantic lexicon. They followed this up with a further study [246] which discussed a number of other deep learning models such as RNN and Bi-LSTM in the domain of Sinhala sentiment analysis. Jayasuriya et al. [278] proposed a method to classify Sinhala posts in the domain of sports into positive and negative class sentiments. Ranathunga and Liyanage [279] claimed that using word embedding models as semantic features can compensate for the lack of well developed language-specific linguistic or language resources in the case of analysing sentiment of Sinhala news comments. Jayasuriya et al. [280] conducted a comparative study between word N-grams and character N-grams in the task of semantic classification of Sinhala content in social media. Which they soon followed up with an ensemble approach [281]. The work by Karunarathne [282] used word embedding to analyse the sentiment of manually annotated Sinhala Tweets. Abeyratne and Jayaratne [283] conducted a multi model analysis on classifying Sinhala songs by emotion.
@@
The work by Jayawickrama et al. [284, 285] used the data set released by Wijeratne and de Silva [99] to predict the reactions induced by Sinhala Facebook posts. They then extended the work [286] and compared the results obtained with their data set against that of Senevirathne et al. [246]Aththanayaka and Naleer [287] used Random forest, Support vector machines, and Multinomial Naïve Bayes models to analyse sentiment in Sinhala-English code-mixed text from social media. The ACTSEA dataset120120120https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al. [135]. The dataset contains 318,308 Sinhala tweets annotated with emotions. The study by Dhananjaya et al. [288] uses a sentiment lexicon of a high‐resource language to fine-tune Pre‐trained multilingual language models (PMLMs) such as mBERT and XLM‐R on an intermediate task which in turn is then used on the sentiment classification task of Sinhala. A sentiment-annotated dataset of Sinhala-English code mixed YouTube comments has been created by Uthpala and Thirukumaran [289]. However, given that even their paper is behind a paywall, there is no free and public way to access this dataset. The work by Gurgurov et al. [230] compared the performance of LLMs for the Sinhala sentiment analysis task. They report the highest results for the Masked Language Modeling objective for Sinhala in both of their configurations where they used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA). Liyanaarachchi et al. [290] used character-level embedding to classify sentiments of Sinhala news comments as positive or negative. They claim that the 1-of-m embedding outperforms Log-m embedding for this task. Bandaranayake and Usoof [291] used BERT, DistilBERT, RoBERTa, and XLM-R to analyse the sentiment of Sinhala news comments [246]. They report that XLMR-large model achieves the best result. Fernando and Ranathunga [292] in their study on improving cross-lingual representation of multilingual language models for low-resource languages using linguistic entity masking used Sinhala code mixed sentiment analysis as one of the evaluations. The work by Peiris and Sumanathilaka [293] analyses emotions in Sinhala songs both in the directions of lyrics and audio. In their analysis of lyrics, they use RoBERTa to represent the Sinhala words. Rizvi et al. [294] used XLM-RoBERTa to extract keywords and classify aspects in Sinhala Code-Mixed data. Then in their follow up work [295] they again used XLM-RoBERTa, this time for sentiment analysis and, on top of that used SHAP [296] and LIME [240] to make the model explainable.
@@ 3.12 Hate Speech Detection
A machine learning approach to detect hate speech in Sinhala was proposed by De Silva [298]. A feature model and a data set121121121https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [128]Sandaruwan et al. [299] have attempted to identify abusive Sinhala comments in social media using text mining and machine learning techniques. A cyberbullying comment classification study for Sinhala was conducted by [300] where they used classical machine learning algorithms. The study by Hettiarachchi et al. [301] used classical machine learning methods to detect hate speech in Romanized Sinhala social media posts. While the basic idea is the same, they have avoided mentioning transliteration in their paper. The study by Samarasinghe et al. [302] proposed using CNNs for detecting hate speech in Sinhala text. Kariyawasam [303] proposed a machine learning approach for identifying toxic Sinhala language on social media. Guruge et al. [304] used an ensemble of Naïve Bayes, Support Vector Machine, XGBoost, MLP, and AdaBoost to detect hate speech in 49019 Tweets they collected from February of 2021 to April of 2021. Sandaruwan et al. [305] collected 3000 Sinhala comments corpus and conducted Multinomial Naïve Bayes to detect hate speech. Sheran [306] claims to have used machine learning to detect hate speech written in Sinhala or Singlish on social media. A deep learning approach for detecting hate speech in Sinhala tweets was explored by Munasinghe and Thayasivam [307]. The study by Shalinda and Munasinghe [308] utilized classical machine learning techniques such as Linear Support vector classifier, Random Forest Classification, SGD classifier, Logistic Regression, XGBoost classifier and multinomial Naive Bayes classifier on both Sinhala and Singlish (Romanized Sinhala) to identify hate words. They report that the SGD classifier over TF-IDF with uni-grams and bi-grams gives them the highest accuracy. Gamage et al. [309] conducted a comparative analysis on a number of embedding systems as well as classical frequency-based methods for Sinhala hate speech detection. The study by Fernando et al. [310] also claims to use machine learning and deep learning to detect hate speech in Sinhala. The study by Perera et al. [311] predicted Sinhala hate speech using user behaviour on twitter by applying ensemble Machine Learning to classify them by the five big personality traits. The follow-up study by Perera et al. [129] analyses Sinhala hate speech propagation on Twitter. They have also published a dataset122122122https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets. Rajapaksha et al. [312] used deep learning to identify trending periods of hate topics on Twitter in Sinhala. Arachchi et al. [313] used a web-based tool to translate between Sinhala and English to detect inappropriate word usage in Sinhala. The study by Ehelepola [314] analyzed Sinhala text from social media and e-commerce sites for hate speech. The work by Dikwatta et al. [315] used supervised algorithms to detect Sinhala hate speech in text from image posts on the Internet (memes). Wickramaarachchi et al. [316] used LSTM on BART to compare the title and description of Sinhala youtube videos to their audio in order to determine if they contain hate speech. The study by Muthuthanthri and Smith [317] used an annotated dataset from Facebook with TF-IDF and fastText embedding features with SVM, CNN, RNN, LSTM as well as BERT and GPT2. They concluded that BERT obtains the best result for detecting hate speech in English-Sinhala code-mixed data. Chavinda and Thayasivam [172] have used Multilingual Large Language Models (MLLMs) with Dual Contrastive Learning (DCL) to detect Sinhala hate speech on Facebook123123123https://huggingface.co/datasets/krishan-CSE/Facebook_Sinhala_Hate_Speech and Twitter124124124https://huggingface.co/datasets/krishan-CSE/Twitter_Sinhala_Hate_Speech data sets. They report TwHIN-BERT-base [318] model to have obtained the best results. The work by Jahnavi and Chaturvedi [319] used BERT with CNN and LSTM to detect hate speech in Sinhala. Perera et al. [320] uses XLM-RoBERTa to classify Sinhala youtube video titles and comments to severity levels of hate. This, by definition is a multi-class classification. The work by Tanjim [321] analyses multiple low-resourced languages in the context of Cyberbullying detection. Ultimately, for the case of Sinhala as the target language, they recommend using Assamese, Bengali, Nepali, Urdu, Hindi, or Bodo as the source language for transfer learning this task.
, title=Multilingual Cyberbullying Detection in Low-Resource Languages and Dialects
@@
After the release of the SOLD dataset by Ranasinghe et al. [140] for the hate speech detection task in Sinhala, a number of subsequent works have been conducted using it. Fernando and Deng [322] used traditional machine learning algorithms to predict Sinhala hate speech. Ranasinghe and Zampieri [323] compared the results of pre-trained mT5 on Sinhala tweets from the SOLD dataset against results for German, Spanish, Hindi, Korean, and Marathi. The work by Bestgen [324] compared the result of classical machine learning algorithms on simple n-gram features for the hate speech detection task on the SOLD dataset for Sinhala against results for Assamese and Bengali. Dmonte et al. [325] used the SOLD data set to determine whether machine translation impacts the performance of hate speech detection. Haturusinghe et al. [326] introduces a series of models named Subasa that are tested on the SOLD dataset. Their trained PMLs125125125https://github.com/haturusinghe/subasa-plm and LLMs126126126https://github.com/haturusinghe/subasa-llm are separately publicly available. Ou6t of these models, they report that their Subasa-XLM-R model outperforms even the state-of-the-art commercial LLMs when evaluated on the SOLD data set. Ranasinghe et al. [327] used the SOLD dataset for the Hate Speech And Offensive Content (HASOC) identification task127127127https://hasocfire.github.io/hasoc/2023/task1.html they introduced. In completing the task, the study by Narayan et al. [328] compared the effectiveness of a number of pre-trained deep learning models for Sinhala hate speech detection. They showed that XLM Roberta Base achieves a macro-f1 of 83.49% over the LSTM baseline of 75.30%. For the same task, Ojo et al. [329] used mBERT to identify hate and offensive content in Sinhala. The work by Rostamkhani and Eetemadi [330] compared the zero-shot performance of ChatGPT on the SOLD dataset against a number of other Sinhala models. They conclude that Zero-shot ChatGPT perform better in Sinhala hate speech detection than models that were not fine-tuned on SOLD. Dmonte et al. [141] used the SOLD data set to verify the new MT-Offense128128128https://github.com/LanguageTechnologyLab/MT-Offense data set they created by using NMT to translate the English offensive language benchmark dataset, OLID [142]. They report the best F1 of 0.550 for the translation done using nllb-200-3.3B [134].